Skip to Search

 
diagonal spacer

Systems Staff Meeting Notes
September 28, 2007

Present: Thomas Bielicke, Cynthia Bowker, Will Chaney, Logan Clapp, Debra Fetzer, Winston Harris, Bobbie Meng, Tom Minton, Chris Nicolich, Stephen Williams
Announcements
  1. Cynthia is going to start decorating for Halloween this weekend.
  2. Talk of: holidays, Winky Dink and 'plucking your magic twanger'.
  3. Library Council Meeting highlights:
    Budget & Hiring Freeze status - the budget is not firmed up yet by Tallahassee. As of right now, they are asking for a 4.2% callback. There should not be a big change once the budget is approved. We may be able to hire vacated positions at that time.
    Staff Marketing Equity - second half of pay raise - the total amount has been approved by Tigert. Best guess would be January - but, it is all contingent on budget approval.
    Student hiring - pushing for Federal Work Study hires.
    Shared remote storage facility - to be located out at ALF. The proposal has been submitted, not yet approved.
    Travel requirements - new form: professional development or official business.
    Employee Recognition Committee - recommends that service pins are handed out as soon as they are received - at the closest possible departmental gathering.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. Laptop batteries - Access Services - Scott brought over five batteries that were dead and we replaced them.
Liaisons
  1. Inventory problems - we're moving towards getting them straightened out.
  2. Verify Patron - computers in West were updated. Room 212 did not have libpublic in the admin group. There is a list of problems for Evan to fix. Reapplied group policy.
  3. Dual monitors on public machine - a patron reported a vulnerability (via snail mail) on the high-end machines in West. Not so much of an issue - they have Deep Freeze.
Networking
  1. McAfee - auto-update is still not functioning. Must update manually.
  2. SAN - coming along - will power it on today. The switches and SAN are configured. Trying to figure out the routing. Should have it up and running by next week, then will be able to demo VMware.
  3. Wireless in East - AP's to be installed.
  4. Backup Exec jobs failed last night.
Software Projects Status
  1. Blog & Wiki server - is installed. Laurie and Missy have blogs. Next, we'll add the rest of the Web 2.0 group. Matthew sent an email asking about the security. IIS is the easiest to use. Logan found info on LDAP authentication - two solutions to investigate.
  2. Grover - working on the tickets page. No estimate for completion date yet.
  3. Tracking database - version 3.93 has been tested, approved & is in production.
  4. VIVO - Valrie sent an email requesting a letter of support for the mini-grant committee. It shouldn't be a problem to run; we can put it on our server.
  5. Facilities tickets - changes completed.
  6. Problem with Patron Update - a patron kept getting expiration messages. It had to do with his fellowship status. Winston made changes to fix.
  7. Logan set up new disk drives for the JPEG2K server. We are in the process of copying files.
Web Projects
  1. Tom has made all of the new phone number changes in the Staff Directory. The only 392's left are tape recordings and faxes.
  2. The virtual server is working - server-side includes. Blake has converted. Tom is waiting for a policy change. He will conduct a class on server-side includes.
  3. Fundraising site - Tom as a meeting with Barbara Hood on Tuesday. Will start working on it next week.
Systems Security Management
  1. Meeting held.

 

University of Florida Home Page